# What makes an image realistic?

### v3

@@ 4.2 Universal tests
where $\eta$ is a parameter which controls the trade-off between false positive and false negatives. Note that the left-hand side is our universal critic. If we accept the Neyman-Pearson lemma then it is easy to accept that our measure of realism should take the form of a likelihood ratio instead of just $P({\mathbf{x}})$. However, this does not yet explain why our choice of alternative hypothesis should be $S$.
@@
for all ${\mathbf{x}}^{B}$ and all $B$ (following the same reasoning as in Eqs. 17-20). That is, $U^{B}$ additively dominates any computable likelihood ratio test and the constant $\log(1/\pi_{Q})$ becomes negligible for sufficiently large $B$. Asymptotically, the universal critic is as sensitive to unrealistic examples as any other test based on an alternative hypothesis $Q$ [^4] Li and Vitányi (1997, Chapter 4.3) proved the stronger result that randomness deficiency additively dominates any so-called sum-$P$ test.
@@ 4.3 MCMC
where $P(n\mid\mathbf{x}_{t})\propto\pi_{n}q_{n}({\mathbf{x}}_{t})$ (Appendix B). Note the resemblance of the right-hand side to our universal critic. If we restrict the universal critic to distributions with differentiable densities, then gradient descent on its density can be viewed as a Bayesian’s attempt to simulate Eq. 24.
@@ 5 Related work
Given the wide range of related fields and the vast amount of work in them (Section 1), it is impossible to review any meaningful fraction of related work here. Instead, we will focus on two successful examples with interesting but hitherto unexplored connections to universal critics.
@@ 5.2 Score distillation sampling
where $w(t)$ are hyperparameters assignings weights to the different noise levels. Is $\mathcal{L}_{\text{SDS}}$ a good candidate for $U$? We can see that SDS tries to find ${\mathbf{x}}$ such that ${\mathbf{z}}_{t}$ is near modes of $p_{t}$. Note that $p_{t}$ is essentially the density of ${\mathbf{x}}$ smoothed via convolution with a Gaussian kernel, and so SDS appears fundamentally similar to using $p$ as a measure of realism (Section 2.1) and susceptible to similar failures. Indeed, if the data distribution itself is Gaussian, then $p_{t}$ is Gaussian and the optimal ${\mathbf{x}}$ becomes the mean, which tends to be unrealistic. This raises the question of why SDS performs well in practice. The key to its success lies in classifier-free guidance (CFG; Ho and Salimans, 2021). Instead of using $\hat{\epsilon}_{t}$ directly, this now common trick uses
$\hat{\epsilon}_{t}^{v}({\mathbf{z}}_{t};y)=(1+v)\hat{\epsilon}_{t}({\mathbf{z}}_{t};y)-v\hat{\epsilon}_{t}({\mathbf{z}}_{t})$ (30)
instead, where $\hat{\epsilon}_{t}({\mathbf{z}}_{t})$ is an unconditional prediction of $\bm{\epsilon}$ and the guidance weight $v\geq 0$ is a hyperparemeter. This corresponds to a gradient signal proportional to
@@
We therefore submit that the reason SDS works well is not explained by its ability to find modes in densities or its connections to model distillation, but by its ability to approximate universal critics. Reinterpreting SDS in this way suggests new ways of overcoming its weaknesses (e.g., its tendency to produce oversaturated images), such as a more intentional design of the mixture of alternatives, or batched losses analogous to Eq. 16.
@@ 6 Discussion
In this position paper we have argued that the question of realism is equivalent to the question of whether observations originated from a particular distribution. This allowed us to draw on insights from algorithmic information theory and to propose universal critics, or randomness deficiency (Li and Vitányi, 1997), as a rational answer. Perceptual quality can be seen as the result of a (necessarily) imperfect approximation of universal critics. However, despite the relevance of these concepts to realism and related problems in machine learning, discussions of randomness deficiency are surprisingly absent from its literature. Instead, dominant notions of realism continue to be based on probability (e.g., Ruff et al., 2021; Poole et al., 2023), typicality (e.g., Nalisnick et al., 2019b) or divergences (e.g., Blau and Michaeli, 2018; Theis and Wagner, 2021). An important open question for future research is how to make universal critics practical, through implicit or explicit approximations of Eqs. 8 and 16.
@@ A.2 Strong typicality
Finally, we are unaware of generalizations of strong typicality to continuous spaces, making it less applicable in settings where we want to optimize for realism. That said, maximum mean discrepancy [MMD; Gretton et al., 2012] is similar in spirit and can be applied in continuous spaces. We discuss MMD in Section 3.2.
@@ Appendix C Limited-memory observer
Here we elaborate on the relationship between the batched universal critic and an ideal observer in a sequential prediction task. The observer assigns values $T({\mathbf{x}})$ to images ${\mathbf{x}}$. Assume we ask the observer to maximize
@@
that is, the observer receives a reward of $T({\mathbf{x}})$ if ${\mathbf{x}}\sim Q$ and a penalty of $\exp(T({\mathbf{x}}))$ if ${\mathbf{x}}\sim P$. The optimal critic is given by [Glaser et al., 2021]

### v4

@@ 4.2 Universal Tests
where $\eta$ is a parameter which controls the trade-off between false positives and false negatives. Note that the left-hand side is our universal critic. If we accept the Neyman-Pearson lemma then it is easy to accept that our measure of realism should take the form of a likelihood ratio instead of just $P({\mathbf{x}})$. However, this does not yet explain why our choice of alternative hypothesis should be $S$.
@@
for all ${\mathbf{x}}^{B}$ and all $B$ (following the same reasoning as in Eqs. 20-23). That is, $U^{B}$ additively dominates any computable likelihood ratio test and the constant $\log(1/\pi_{Q})$ becomes negligible for sufficiently large $B$. Asymptotically, the universal critic is as sensitive to unrealistic examples as any other test based on an alternative hypothesis $Q$ [^6] Li and Vitányi (1997, Chapter 4.3) proved the stronger result that randomness deficiency additively dominates any so-called sum-$P$ test.
@@ 4.3 MCMC
where $P(n\mid\mathbf{x}_{t})\propto\pi_{n}q_{n}({\mathbf{x}}_{t})$ (Appendix B). Note the resemblance of the right-hand side to Solomonoff’s probability. If we restrict the universal critic to distributions with differentiable densities, then gradient descent on its density can be viewed as a Bayesian’s attempt to simulate Eq. 27.
@@ 5 Related Work
Given the wide range of related fields and the vast amount of work in them (Section 1), it is impossible to review any meaningful fraction of related work here. Instead, we will focus on two successful examples with interesting connections to universal critics.
@@ 5.1 Input Complexity
We note that input complexity has also been considered in statistics for its applications in hypothesis testing, including as an approximation to universal tests (Ryabko et al., 2006).
@@ 5.2 Score Distillation Sampling
where $w(t)$ are hyperparameters assigning weights to the different noise levels. Is $\mathcal{L}_{\text{SDS}}$ a good candidate for $U$? We can see that SDS tries to find ${\mathbf{x}}$ such that ${\mathbf{z}}_{t}$ is near modes of $p_{t}$. Note that $p_{t}$ is essentially the density of ${\mathbf{x}}$ smoothed via convolution with a Gaussian kernel, and so SDS appears fundamentally similar to using $p$ as a measure of realism (Section 2.1) and susceptible to similar failures. Indeed, if the data distribution was Gaussian, then $p_{t}$ would also be Gaussian and the optimal ${\mathbf{x}}$ would be the mean, which tends to be unrealistic. This raises the question of why SDS performs well in practice. The key to its success lies in classifier-free guidance (CFG; Ho and Salimans, 2021). Instead of using $\hat{\epsilon}_{t}$ directly, this common trick is to use
$\hat{\epsilon}_{t}^{v}({\mathbf{z}}_{t};y)=(1+v)\hat{\epsilon}_{t}({\mathbf{z}}_{t};y)-v\hat{\epsilon}_{t}({\mathbf{z}}_{t}),$ (33)
where $\hat{\epsilon}_{t}({\mathbf{z}}_{t})$ is an unconditional prediction of $\bm{\epsilon}$ and the guidance weight $v\geq 0$ is a hyperparemeter. This corresponds to a gradient signal proportional to
@@
We therefore submit that the reason SDS works well is not explained by its ability to find modes in densities or its connections to model distillation techniques, but by its ability to approximate universal critics. Reinterpreting SDS in this way suggests new ways of overcoming its weaknesses (e.g., its tendency to produce oversaturated images), such as a more intentional design of the mixture of alternatives, or batched losses analogous to Eq. 19.
@@ 6 Discussion
In this position paper we have argued that the question of realism is equivalent to the question of randomness, that is, whether observations originated from a particular distribution. This allowed us to draw on insights from algorithmic information theory and to propose universal critics, or randomness deficiency (Li and Vitányi, 1997), as a rational answer. Perceptual quality can be viewed as the result of a (necessarily) imperfect approximation of universal critics. However, despite the relevance of these concepts to problems in machine learning, discussions of randomness deficiency are notably absent from its literature. Instead, dominant notions of realism continue to be based on probability (e.g., Ruff et al., 2021; Poole et al., 2023), typicality (e.g., Nalisnick et al., 2019b) or divergences (e.g., Blau and Michaeli, 2018; Theis and Wagner, 2021).
A divergence of zero is a sufficient condition for perfect realism but corresponds to an ideal observer with access to an infinite stream of examples. As such, it is stronger than required for most practical applications where observers only have access to one or a few examples. At the other end of the spectrum, weak typicality is an example of a criterion which only considers a necessary criterion, while most no-reference metrics correspond to neither a necessary nor a sufficient criterion (e.g., high probability in some feature space). Universal critics enable principled relaxations of divergence-based constraints. While weaker than divergences (in the desired way), they are still strong in the sense that they are as strong as other likelihood-ratio tests for realism, up to a constant which depends on the complexity of the competing test (Section 4.2).
Many interesting practical and theoretical questions remain. For example, what is the impact of different choices of $\pi_{n}$ on the sample efficiency of universal critics? What are the implications of using universal critics in rate-distortion-realism trade-offs? Most importantly, what do practical approximations to universal critics (Eqs. 8 or 19) look like that can serve as optimization targets?
@@ A.2 Strong typicality
Here we consider strong typicality. A closely related notion ($P$-typicality) was considered by Chen et al. [2022] to quantify the realism of a batch of examples. Strong typicality is also similar in spirit to maximum mean discrepancy [MMD; Gretton et al., 2012], which was discussed in Section 3.2.
@@ Appendix C Limited-memory observer
Here we elaborate on the relationship between the batched universal critic and an ideal observer in a sequential prediction task. Assume an observer assigns a value $T({\mathbf{x}})$ to an image ${\mathbf{x}}$. Further assume we ask the observer to
@@
that is, the observer receives a reward of $T({\mathbf{x}})$ if ${\mathbf{x}}\sim Q$ and a penalty of $\exp(T({\mathbf{x}}))$ if ${\mathbf{x}}\sim P$. The optimal output is then given by [Glaser et al., 2021]